# Surrogate Neural Networks Local Stability for Aircraft Predictive Maintenance

### v2

@@ 6 Conclusions and ways forward
A Complete verification technique using MILP
A.1 MILP encoding of the neural network
The MILP encoding makes use of bounds computed by symbolic interval propagation. For each layer $i$ and neuron $j$ in the layer $i$, two continuous variables are created : $y_{i,j,-},y_{i,j,+}$ which corresponds to pre-activation and post-activation values of the neurons. Bounds of these variables are known thanks to the bound propagation function. Layer $i=0$ corresponds to input values.
The MILP is encoded as follow:
  1. 1.
  2. 2.
The $max$ constraint is native into the Gurobi library and can be chosen instead of the classically used Big-M constraint (e.g., in Botoeva et al. [2020a]).
A.2 Property encoding
Given $x,f(x),\underline{\delta},\overline{\delta}$ defined in property 1, we encode the negation of the property as follows:
  1. 1.
    Let $\forall j\in[1,81],indic_{-,j},indic_{+,j}$ be binary variables.
  2. 2.
  3. 3.
  4. 4.
    $\sum_{j}{(indic_{+,j}+indic_{-,j})}\geq 1$
Each indicator variable will encode the fact that for one output index, the perturbed NN prediction goes out of the bow-tie property. The last constraint thus encodes that the solver has to find a counter-example violating the property. If the MILP solver finds a solution, we conclude that the property is violated. On the contrary, if the solver proves the infeasibility of the model, we conclude that the property is verified.

### v3

@@ 7 Conclusions and future work
A Open-source verification pipeline
The open-source code of the verification pipeline (AIROBAS) is made available on github.
It provides functionalities to:
  • load input data and models,
  • define a robustness property to be verified. The code has been implemented to be as generic as possible and highly parametrizable regarding property definition,
  • design customized verification pipeline,
  • run verification with a default set of empirical and formal methods e.g., attacks using the cleverhans library, incomplete verification using LiRPA-based functionalities implemented in decomon etc.,
  • append new verification functions or link the code to open-source libraries of one’s choice.
It also provides a set of step-by-step notebooks to train a number of NN surrogate models (e.g., surrogate to the Rosenbrock analytic function, a runway braking distance estimate etc.), define robustness properties to be assessed and proceed to prove or disprove such properties for a set of test data.